# SemiCD-VL: Visual-Language Model Guidance Makes Better Semi-supervised Change Detector

### v4

@@ I Introduction
The rest of this paper is organized as follows. Section 2 briefly introduces supervised CD, semi-supervised learning and dense prediction with VLMs. Section 3 presents the proposed semi-supervised framework and the implementation details. The extensive experimental results and detailed discussion are presented in Section 4. Finally, our conclusions are given in Section 5.
@@ III-A Preliminaries
In Semi-supervised CD, we have a small set of labeled data $\mathcal{D}_{l}=\{\mathbf{x}_{i}^{l},\mathbf{y}_{i}^{l}\}_{i=1}^{N_{l}}$ and a large set of unlabeled data $\mathcal{D}_{u}=\{\mathbf{x}_{i}^{u}\}_{i=1}^{N_{u}}$, where $\mathbf{x}_{i}^{*}=(\mathbf{t_{1}}_{i}^{*},\mathbf{t_{2}}_{i}^{*})$ and $*\in\{l,u\}$, $\mathbf{t_{1}}_{i}^{*}$ and $\mathbf{t_{2}}_{i}^{*}$ denote the pre-event and post-event images, $N_{l}$ and $N_{u}$ denote the number of labeled and unlabeled images and $N_{u}\gg N_{l}$. Following the current mainstream weak-to-strong consistency regularization methods (e.g. FixMatch [20]) in SSL, we train our model on both labeled and unlabeled data. Specifically, the overall loss $\mathcal{L}_{cr}$ under consistency regularization framework consists of two cross-entropy loss terms, i.e., supervised loss $\mathcal{L}_{s}$ and un-supervised loss $\mathcal{L}_{u}$:
@@
$\mathcal{L}_{s}$ is a standard pixel-wise cross-entropy loss, which is calculated on weakly perturbed labeled samples:
@@ III-B1 Category Definition & Single-temporal Reasoning
Figure 3: Visualisation of direct inference using VLM with prompts: house, building, road, grass, tree, water. (The color rendering is random, just to distinguish different categories.) 111powered by: https://huggingface.co/spaces/shenyunhang/APE
@@ III-C VLM Guidance
The pseudo label $I^{\textit{mix-diff}}$ generated by VLM can provide additional beneficial guidance for unlabeled data. Under the FixMatch paradigm, for unlabeled data, there are two types of predictions, $y^{w}$ generated by $x^{u}$ via weak perturbations and $y^{s}$ generated via strong perturbations. To fully utilize the pseudo label $I^{\textit{mix-diff}}$ generated by the VLM, we build supervision from $I^{\textit{mix-diff}}$ for both $y^{w}$ and $y^{s}$ via a cross entropy loss $\mathcal{L}_{vl}$:
@@
We argue that regularizing the weak/strong perturbation predictions with a shared VLM label can also be seen as enforcing the consistency between these two predictions. On the other hand, our $\mathcal{L}_{vl}$ shares the core spirits of contrastive learning [12, 42, 43]. Suppose $(q_{w},q_{s})$ are feature vectors of the weakly/strongly perturbed views of sample $x^{u}$, and $h_{+}$ is the classifier weight of the class matched to $I^{\textit{mix-diff}}$. When adopting cross entropy loss, $q_{j}\cdot h_{+}$ is maximized against $\sum_{a\in A}{q_{j}\cdot h_{a}}$, where $j\in\{w,s\}$, and $h_{a}$ is classifier weight of concept class $a$. This process is also maximizing the similarity between $q_{w}$ and $q_{s}$, which can be approximated as InfoNCE loss [44]:
@@ III-D Dual Projection Head
Then, the hard label $\hat{\mathbf{y}}^{w}$ predicted by the weakly perturbed image is applied to supervise $y^{s}_{cr}$, as formulated in Eq. 3; similarly, the change mask $I^{\textit{mix-diff}}$ generated by the mixed CEG is used to supervise $y^{s}_{vl}$, in the format of cross entropy loss. Furthermore, since $\mathbf{h}_{vl}$ is independent of $\mathbf{h}_{cr}$, we can also construct supervision from $I^{\textit{mix-diff}}$ for the prediction $y^{w}_{vl}$ of weakly perturbed image obtained via $\mathbf{h}_{vl}$, as mentioned in Section III-C.
@@ III-E Decoupled Semantic Guidance
In detail, for decoupling change information, two auxiliary segmentation decoders are constructed to respectively predict the bi-temporal segmentation masks, as shown in Fig. 2 (4). Specifically, these two decoders receive the embeddings of the bi-temporal images from the encoders and up-sample them hierarchically to the high-resolution features $q_{t_{1}}$ and $q_{t_{2}}$. Notably, the two segmentation decoders share the same set of weights, making efficient parameter utilization. Finally, the high-resolution bi-temporal features are fed into a linear classifier $\mathbf{h}_{seg}$ to obtain the segmentation prediction $y_{t_{1}}$ and $y_{t_{2}}$. The segmentation decoders are activated only during training, without introducing extra cost in reasoning. In the training phase, the bi-temporal segmentation decoders enhance and clarifies the feature representation of the encoder through back-propagation, which explicitly employs single-temporal scene interpretation as a prior task of CD and decouples the bi-temporal entangled features.
@@ III-F Contrastive Consistency Regularization
$\mathcal{L}_{c}=\left\{\begin{aligned} \frac{1}{n_{u}}\sum\mathcal{D}\left(q_{t_{1}},q_{t_{2}}\right),\hat{y}=0\\ \frac{1}{n_{c}}\sum\max\left(0,\epsilon-\mathcal{D}\left(q_{t_{1}},q_{t_{2}}\right)\right),\hat{y}=1\end{aligned}\right.,$ (15)
@@ IV-E2 Components of SemiCD-VL
TABLE IX: Ablation study of weights $\lambda_{c}$ and $\lambda_{vl}$ for contrastive loss and VLMguidance loss.
$\lambda_{c}$ $\lambda_{vl}$ Presicion Recall $IoU^{c}$

### v5

@@ I Introduction
The rest of this paper is organized as follows. Section II briefly introduces supervised CD, semi-supervised learning and dense prediction with VLMs. Section III presents the proposed semi-supervised framework and the implementation details. The extensive experimental results and detailed discussion are presented in Section IV. Finally, our conclusions are given in Section V.
@@ III-A Preliminaries
In Semi-supervised CD, we have a small set of labeled data $\mathcal{D}_{l}=\{\mathbf{x}_{i}^{l},\mathbf{y}_{i}^{l}\}_{i=1}^{N_{l}}$ and a large set of unlabeled data $\mathcal{D}_{u}=\{\mathbf{x}_{i}^{u}\}_{i=1}^{N_{u}}$, where $\mathbf{x}_{i}^{*}=(\mathbf{t_{1}}_{i}^{*},\mathbf{t_{2}}_{i}^{*})$ and $*\in\{l,u\}$, $\mathbf{t_{1}}_{i}^{*}$ and $\mathbf{t_{2}}_{i}^{*}$ denote the pre-event and post-event images, $N_{l}$ and $N_{u}$ denote the number of labeled and unlabeled images and $N_{u}\gg N_{l}$. Following the current mainstream weak-to-strong consistency regularization methods (e.g. FixMatch [20]) in SSL, we train our model on both labeled and unlabeled data. Specifically, the overall loss $\mathcal{L}_{cr}$ under consistency regularization framework consists of two cross-entropy (CE) loss terms, i.e., supervised loss $\mathcal{L}_{s}$ and un-supervised loss $\mathcal{L}_{u}$:
@@
$\mathcal{L}_{s}$ is a standard pixel-wise CE loss, which is calculated on weakly perturbed labeled samples:
@@ III-B1 Category Definition & Single-temporal Reasoning
Figure 3: Visualization of direct inference using VLM with prompts: house, building, road, grass, tree, water. (The color rendering is random, just to distinguish different categories.) 111powered by: https://huggingface.co/spaces/shenyunhang/APE
@@ III-C VLM Guidance
The pseudo label $I^{\textit{mix-diff}}$ generated by VLM can provide additional beneficial guidance for unlabeled data. Under the FixMatch paradigm, for unlabeled data, there are two types of predictions, $y^{w}$ generated by $x^{u}$ via weak perturbations and $y^{s}$ generated via strong perturbations. To fully utilize the pseudo label $I^{\textit{mix-diff}}$ generated by the VLM, we build supervision from $I^{\textit{mix-diff}}$ for both $y^{w}$ and $y^{s}$ via a CE loss $\mathcal{L}_{vl}$:
@@
We argue that regularizing the weak/strong perturbation predictions with a shared VLM label can also be seen as enforcing the consistency between these two predictions. On the other hand, our $\mathcal{L}_{vl}$ shares the core spirits of contrastive learning [12, 42, 43]. Suppose $(q_{w},q_{s})$ are feature vectors of the weakly/strongly perturbed views of sample $x^{u}$, and $h_{+}$ is the classifier weight of the class matched to $I^{\textit{mix-diff}}$. When adopting CE loss, $q_{j}\cdot h_{+}$ is maximized against $\sum_{a\in A}{q_{j}\cdot h_{a}}$, where $j\in\{w,s\}$, and $h_{a}$ is classifier weight of concept class $a$. This process is also maximizing the similarity between $q_{w}$ and $q_{s}$, which can be approximated as InfoNCE loss [44]:
@@ III-D Dual Projection Head
Then, the hard label $\hat{\mathbf{y}}^{w}$ predicted by the weakly perturbed image is applied to supervise $y^{s}_{cr}$, as formulated in Eq. 3; similarly, the change mask $I^{\textit{mix-diff}}$ generated by the mixed CEG is used to supervise $y^{s}_{vl}$, in the format of CE loss. Furthermore, since $\mathbf{h}_{vl}$ is independent of $\mathbf{h}_{cr}$, we can also construct supervision from $I^{\textit{mix-diff}}$ for the prediction $y^{w}_{vl}$ of weakly perturbed image obtained via $\mathbf{h}_{vl}$, as mentioned in Section III-C.
@@ III-E Decoupled Semantic Guidance
In detail, for decoupling change information, two auxiliary segmentation decoders are constructed to respectively predict the bi-temporal segmentation masks, as shown in Fig. 2 (4). Specifically, these two decoders receive the embeddings of the bi-temporal images from the encoders and up-sample them hierarchically to the high-resolution features $q_{t_{1}}$ and $q_{t_{2}}$. Notably, the two segmentation decoders share the same set of weights, making efficient parameter utilization. Finally, the high-resolution bi-temporal features are fed into a linear classifier $\mathbf{h}_{seg}$ to obtain the segmentation prediction $y_{t_{1}}$ and $y_{t_{2}}$. These two predictions are supervised by the segmentation masks $I_{1}$ and $I_{2}$ (filtered by $\beta$) provided by the VLM via a CE loss and added as part of $\mathcal{L}_{vl}$. The segmentation decoders are activated only during training, without introducing extra cost in reasoning. In the training phase, the bi-temporal segmentation decoders enhance and clarifies the feature representation of the encoder through back-propagation, which explicitly employs single-temporal scene interpretation as a prior task of CD and decouples the bi-temporal entangled features.
@@ III-F Contrastive Consistency Regularization
$\mathcal{L}_{ct}=\left\{\begin{aligned} \frac{1}{n_{u}}\sum\mathcal{D}\left(q_{t_{1}},q_{t_{2}}\right),\hat{y}=0\\ \frac{1}{n_{c}}\sum\max\left(0,\epsilon-\mathcal{D}\left(q_{t_{1}},q_{t_{2}}\right)\right),\hat{y}=1\end{aligned}\right.,$ (15)
@@ IV-E2 Components of SemiCD-VL
TABLE IX: Ablation study of weights $\lambda_{ct}$ and $\lambda_{vl}$ for contrastive loss and VLMguidance loss.
$\lambda_{ct}$ $\lambda_{vl}$ Presicion Recall $IoU^{c}$